OpenStudy (anonymous):

How much money should be deposited today in an account that earns 9.5% compounded monthly so that it will accumulate to $10,000 in three years? (Do not round until the final answer. Then round to the nearest cent.)

OpenStudy (anonymous):

Alright i'll explain this one...

OpenStudy (anonymous):

im not a retard i swear finals are just making me brain dead haha

OpenStudy (anonymous):

haha no no i know exactly how you feel. im right in the middle of finals too

OpenStudy (anonymous):

FV=PV*(1+r/n)^(nt) where r is the rate, n is the compounding periods per year and t is the time in years. So here we're gonna plug the known values in and solve for the unknown value of PV: $10,000=PV*(1+(.095/12))^(12*3) $10,000=PV*(1.00792))^(12*3) $10,000=PV*1.32827 PV=10,000/1.32827=$7,528.59

OpenStudy (anonymous):

yesssssss i got the right answer

OpenStudy (anonymous):

haha nice